OpenStudy (anonymous):

y varies inversely with the square of x, y also varies with x+1 direclty.What is the ratio of the constant \[\frac{k_1}{k_2}\] if x=1

OpenStudy (anonymous):

\[\frac{k_1}{k_2}\]

OpenStudy (anonymous):

i have been looking at alot of proportion problems on OS heres my own

hero (hero):

It's easier than you think

hero (hero):

\(y = \frac{k_1}{x^2}\) \(y = k_2(x + 1)\) x = 1 \(y = \frac{k_1}{1^2} = \frac{k_1}{1}\) \(y + k_2(1 + 1) = 2k_2\) y = y \(\frac{k_1}{1} = 2k_2\) \(\frac{k_1}{k_2} = 2\)

hero (hero):

@Jonask

OpenStudy (anonymous):

you win

hero (hero):

lol, okay :D

OpenStudy (anonymous):

good training for such questions ....